OpenStudy (anonymous):

How would I solve 2|x-3|+1=7?

OpenStudy (anonymous):

Think of the absolute value being parentheses so 2(x-3)+1=7

OpenStudy (anonymous):

Yep Hope it helps Good Luck :)

OpenStudy (wolfe8):

Doesn't the absolute value mean it will give you x-3=3 and x-3=-3 ?